OpenStudy (anonymous):

How do I do this problem? 7+3 * [ 8+8 *( 5+9)] they say the answer is 367 but I can't get it right? Thanx

OpenStudy (anonymous):

I agree with "they"

OpenStudy (anonymous):

Work from the inner most bracket out!

OpenStudy (anonymous):

@rgarcia876 Can you start???

OpenStudy (anonymous):

Yeah but the answer doesn't come out right for me. This is how I've done it 5+9 = 14 = 7+3 * [8+8 * 14] then I do 8+8 = 16 7+3 * [16 *14] then 16*14 = 224 7 + 3 * 224 and then i get stuck?! Can someone explain a little how they do it?

OpenStudy (anonymous):

Nope... do 8 times 14 after 5+9... PEMDAS within each set of parenthesis

OpenStudy (anikate):

do inner brackets then multiply by 8. then add 8. then times that by 3. then add 7.

OpenStudy (anonymous):